A man wishes to have an amount of P2,769.84 after 5 years. He deposited P500 each year into a savings bank that pay “r” percent per annum compounded continuously. Compute the value of “r”.

1
Expert's answer
2022-01-21T09:16:42-0500

This is an annuity

A=500

FV=2769.84

"2769.84=500(\frac{(1+r)^5-1}{r})"


"5.53968=(\frac{(1+r)^5-1}{r})"

"r^5-5.53968r=0"

"r(r^4-5.53968)=0"

r1=0

r2=1.53